# Configuration spaces of circles in the plane

### v1

@@ 3.3. Connected components
Let $T$ and $U$ be labeled trees with $n$ non-root vertices. Then $\varkappa_{T}$ and $\varkappa_{U}$ lie in the same connected component of $\operatorname{LConf}_{n}(S^{1},\mathbb{R}^{2})$ if and only if $T\cong U$. Moreover, every connected component of $\operatorname{LConf}_{n}(S^{1},\mathbb{R}^{2})$ contains some $\varkappa_{T}$ for $T$ a labeled tree, and $T\cong T_{\kappa}$ for each $\kappa$ in this connected component.
@@
Suppose $\varkappa_{T}$ and $\varkappa_{U}$ lie in the same connected component, and we want to show that $T\cong U$. Recall that two labeled trees with the same number of vertices are isomorphic if and only if for all $1\leq i\leq n$, the set of labels of children of the vertex labeled $i$ are the same in $T$ and $U$, with an analogous statement for the root. Choose a path from $\varkappa_{T}$ to $\varkappa_{U}$, call it $\gamma(t)$ ($0\leq t\leq 1$). We can write
@@
so $C_{i}(0)=\mathfrak{C}_{i}^{T}$ and $C_{i}(1)=\mathfrak{C}_{i}^{U}$ for each $1\leq i\leq n$. Note that $\gamma(t)$ is a configuration at every time $t$, so if the $j$th circle is nested in the $i$th circle at some time $t$, this must hold for all times $t$. This shows that for any $i$ and $j$, we have that $\mathfrak{C}_{i}^{T}\vdash\mathfrak{C}_{j}^{T}$ if and only if $\mathfrak{C}_{i}^{U}\vdash\mathfrak{C}_{j}^{U}$. Hence, $i\prec j$ in $T$ if and only if $i\prec j$ in $U$ (with an analogous statement about the root), and so indeed $T\cong U$.
Now suppose $T\cong U$, and we need to prove there is a path in $\operatorname{LConf}_{n}(S^{1},\mathbb{R}^{2})$ from $\varkappa_{T}$ to $\varkappa_{U}$. If $n=0$ there is nothing to prove, so assume $n\geq 1$. In particular the root is not a leaf. Fix a label $i$ such that the vertex labeled $i$ in $T$, and hence in $U$, is a leaf. Let $\widehat{T}$ be the labeled tree obtained by removing the leaf labeled $i$ from $T$, and relabeling the remaining non-root vertices by replacing $j$ with $j-1$ for all $j>i$. Similarly define $\widehat{U}$. Note that $\widehat{T}\cong\widehat{U}$, so by induction there is a path in $\operatorname{LConf}_{n-1}(S^{1},\mathbb{R}^{2})$ from $\varkappa_{\widehat{T}}$ to $\varkappa_{\widehat{U}}$. Let us write
@@
for this path, with $0\leq t\leq 1$, so $\widehat{\gamma}(0)=\varkappa_{\widehat{T}}$ and $\widehat{\gamma}(1)=\varkappa_{\widehat{U}}$, or said another way, $C_{j}(0)=\mathfrak{C}_{j}^{T}$ and $C_{j}(1)=\mathfrak{C}_{j}^{U}$ for all $j\neq i$. Note that for each $t$, the “snapshot” at time $t$ is a configuration, i.e., for each $t$ the $C_{j}(t)$ are pairwise disjoint for all $j$. Now choose (for our fixed $i$) an arbitrary family of circles $C_{i}^{\prime}(t)$ ($0\leq t\leq 1$) with initial circle $C_{i}^{\prime}(0)=\mathfrak{C}_{i}^{T}$, such that $(C_{1}(t),\dots,C_{i-1}(t),C_{i}^{\prime}(t),C_{i+1}(t),\dots,C_{n}(t))$ is a configuration for each $t$; intuitively, we are forced to choose $C_{i}^{\prime}(0)=\mathfrak{C}_{i}^{T}$, but otherwise $C_{i}^{\prime}(t)$ just “goes along for the ride” and ends up at some final circle, call it $C_{i}^{\prime}\coloneqq C_{i}^{\prime}(1)$. Write
@@
so $\gamma^{\prime}(t)$ is a path in $\operatorname{LConf}_{n}(S^{1},\mathbb{R}^{2})$ from $\gamma^{\prime}(0)=\varkappa_{T}$ to the configuration
$\kappa^{\prime}\coloneqq(\mathfrak{C}_{1}^{U},\dots,\mathfrak{C}_{i-1}^{U},C_{i}^{\prime},\mathfrak{C}_{i+1}^{U},\dots,\mathfrak{C}_{n}^{U})\text{.}$
Now let $\gamma(t)$ be a path obtained by concatenating $\gamma^{\prime}(t)$ with some path that leaves all but the $i$th circle fixed, and takes the $i$th circle from $C_{i}^{\prime}$ to $\mathfrak{C}_{i}^{U}$. This is possible since none of the circles in $\varkappa_{U}$ are nested in either of $C_{i}^{\prime}$ or $\mathfrak{C}_{i}^{U}$ (since $i$ is a leaf), and a circle in $\varkappa_{U}$ has $C_{i}^{\prime}$ nested in it if and only if it has $\mathfrak{C}_{i}^{U}$ nested in it (since $\gamma^{\prime}(t)$ is a path and $T\cong U$). All in all, we end up with a path $\gamma(t)$ in $\operatorname{LConf}_{n}(S^{1},\mathbb{R}^{2})$ from $\gamma(0)=\varkappa_{T}$ to $\gamma(1)=\varkappa_{U}$.
Finally, we must prove the last claim. Let $\kappa$ be an arbitrary labeled configuration in $\operatorname{LConf}_{n}(S^{1},\mathbb{R}^{2})$, and it suffices to prove that there is a path from $\kappa$ to $\varkappa_{T}$, where $T\coloneqq T_{\kappa}$. In fact, this works in the same way as the previous part. By induction there is a path between the configurations with the $i$th circles deleted, and then we can extend this to a path between the full configurations since the $i$th circle in $\kappa$ and the $i$th circle in $\varkappa_{T}$ have the same nesting properties with respect to the other circles in their respective configurations. ∎
@@
Let $T$ and $U$ be trees with $n$ non-root vertices. Then $\varkappa_{T}$ and $\varkappa_{U}$ lie in the same connected component of $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$ if and only if $T\cong U$. Moreover, every connected component of $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$ contains some $\varkappa_{T}$ for $T$ a tree, and $T\cong T_{\kappa}$ for each $\kappa$ in this connected component.
@@
First suppose $\varkappa_{T}$ and $\varkappa_{U}$ lie in the same connected component, and we need to show that $T\cong U$. Choose some labeling of $T$, and write $\widetilde{T}$ for the resulting labeled tree, so $p(\varkappa_{\widetilde{T}})=\varkappa_{T}$. We can lift a path in $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$ from $\varkappa_{T}$ to $\varkappa_{U}$, to a path in $\operatorname{LConf}_{n}(S^{1},\mathbb{R}^{2})$ starting at $\varkappa_{\widetilde{T}}$. The end of this path is some labeled configuration that maps under $p$ to $\varkappa_{U}$, and so equals $\varkappa_{\widetilde{U}}$ for some labeled tree $\widetilde{U}$ with underlying tree $U$. By Proposition 3.7, the fact that $\varkappa_{\widetilde{T}}$ and $\varkappa_{\widetilde{U}}$ share a connected component implies that $\widetilde{T}\cong\widetilde{U}$, so in particular $T\cong U$.
Now suppose $T\cong U$. Choose a labeling of $T$, and write $\widetilde{T}$ for the resulting labeled tree. Carry the labels on the vertices of $T$ through a choice of isomorphism $T\to U$, to obtain labels on the vertices of $U$, so the resulting labeled tree $\widetilde{U}$ is isomorphic to $\widetilde{T}$ as a labeled tree. By Proposition 3.7, $\varkappa_{\widetilde{T}}$ and $\varkappa_{\widetilde{U}}$ share a connected component in $\operatorname{LConf}_{n}(S^{1},\mathbb{R}^{2})$. Now the image of a path from $\varkappa_{\widetilde{T}}$ to $\varkappa_{\widetilde{U}}$ under the map $p$ is a path in $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$ from $\varkappa_{T}$ to $\varkappa_{U}$.
@@
For $T$ a labeled tree with $n$ non-root vertices, write $\operatorname{LConf}_{T}(S^{1},\mathbb{R}^{2})$ for the connected component of $\operatorname{LConf}_{n}(S^{1},\mathbb{R}^{2})$ containing $\varkappa_{T}$, and call $\operatorname{LConf}_{T}(S^{1},\mathbb{R}^{2})$ the space of labeled $T$-configurations of $n$ circles in the plane. Similarly, for $T$ an (unlabeled) tree with $n$ non-root vertices, write $\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2})$ for the connected component of $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$ containing $\varkappa_{T}$, and call $\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2})$ the space of $T$-configurations of $n$ circles in the plane.
@@
Figure 12. Up to isomorphism there are exactly two (unlabeled) trees with $2$ non-root vertices, pictured here. These correspond to (unlabeled) configurations of $2$ circles in the plane; the first corresponds to the case where the circles are not nested, and the second to the case where the circles are nested.
@@ 4. Braiding automorphism groups of finite trees
Here the proposition is phrased using the “external” viewpoint of semidirect products, i.e., $G=N\rtimes H$ means that $N$ and $H$ are groups, and $H$ acts on $N$ by automorphisms via some implicitly understood (left) action, so the group operation on $N\rtimes H$ is given by $(n,h)(n^{\prime},h^{\prime})=(n(h.n^{\prime}),hh^{\prime})$. In the proof we will actually decompose $\operatorname{Aut}(T)$ using the “internal” viewpoint, i.e., $G=N\rtimes H$ means that $N$ and $H$ are subgroups of $G$, $N$ is normal, $N\cap H=\{1\}$, and $G=NH$.
@@
Now to prove the decomposition, we will use the internal semidirect product viewpoint. By Lemma 4.1 and its proof, $i$ and $j$ share a block of $\Pi$ if and only if there is an automorphism of $T$ that swaps $T(v)$ and $T(w)$ and fixes everything outside $T(v)\cup T(w)$. These automorphisms generate a copy of $S_{m}^{\Pi}$ inside $\operatorname{Aut}(T)$, since within each block $A\in\Pi$, they generate the relevant copy of $S_{|A|}$. Write $H$ for this subgroup of $\operatorname{Aut}(T)$ isomorphic to $S_{m}^{\Pi}$. Write $N$ for the pointwise fixer in $\operatorname{Aut}(T)$ of $\{v_{1},\dots,v_{m}\}$, so $N$ is isomorphic to $\operatorname{Aut}(T(v_{1}))\times\cdots\times\operatorname{Aut}(T(v_{m}))$. Now observe that every element of $\operatorname{Aut}(T)$ can be written as an automorphism in $H$ composed with an automorphism in $N$, so $\operatorname{Aut}(T)=NH$. Moreover, since $S_{m}^{\Pi}$ acts faithfully on $\{v_{1},\dots,v_{m}\}$, we have $N\cap H=\{1\}$. Finally, $N$ is normal, being the kernel of the action on $\{v_{1},\dots,v_{m}\}$, so we conclude that $\operatorname{Aut}(T)=N\rtimes H$. The conjugation action of $H$ on $N$ coincides with the action of $S_{m}^{\Pi}$ on $\operatorname{Aut}(T(v_{1}))\times\cdots\times\operatorname{Aut}(T(v_{m}))$ by permuting entries, so we are done. ∎
@@
This recursive definition makes sense since each $T(v_{i})$ has strictly fewer vertices than $T$. We get that $\operatorname{BAut}(T)$ is built out of iterated semidirect products of braid groups. To be clear, in the bottom case when $T$ is the trivial tree, so $m=0$, we have that $B_{0}$ is the trivial group (being generated by the empty set), so for $T$ the trivial tree we have that $\operatorname{BAut}(T)$ is the trivial group.
@@
As for the pure case, it is clear from Definition 4.5 that $\operatorname{PBAut}(T)\cong PB_{3}\times PB_{3}\times PB_{2}$ for this $T$, since two vertices have $3$ children, one vertex has $2$ children, and all other vertices have $1$ or 00 children.
@@
If every non-root vertex of $T$ is a child of the root, and there are $n$ non-root vertices, then $\pi_{1}(\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2}))\cong\operatorname{BAut}(T)$ by virtue of both being isomorphic to $B_{n}$.
@@
First note that $\operatorname{BAut}(T)\cong B_{n}$ in this case by construction. Now for $C$ a circle in the plane, let $D_{C}$ be the disk in the plane obtained by taking the union of $C$ with the bounded component of its complement. Since every non-root vertex of $T$ is a child of the root, every configuration $\{C_{1},\dots,C_{n}\}$ in $\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2})$ has the property that no $C_{j}$ is nested in any $C_{i}$. Thus, we have a well defined map $\{C_{1},\dots,C_{n}\}\mapsto\{D_{C_{1}},\dots,D_{C_{n}}\}$ from $\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2})$ to $\operatorname{Conf}_{n}(D^{2},\mathbb{R}^{2})$. This map is clearly a homeomorphism, so $\pi_{1}(\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2}))\cong B_{n}$ by Lemma 2.6. ∎
An analogous procedure shows that $\pi_{1}(\operatorname{LConf}_{T}(S^{1},\mathbb{R}^{2}))\cong\operatorname{PBAut}(T)$ for trees $T$ of this form, thanks to both being isomorphic to $PB_{n}$.

### v2

@@ 3.3. Connected components
Let $T$ and $U$ be labeled trees with $n$ non-root vertices. Then $\underline{\kappa}_{T}$ and $\underline{\kappa}_{U}$ lie in the same connected component of $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$ if and only if $T\cong U$. Moreover, every connected component of $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$ contains some $\underline{\kappa}_{T}$ for $T$ a labeled tree, and $T\cong T_{\kappa}$ for each $\kappa$ in this connected component.
@@
Suppose $\underline{\kappa}_{T}$ and $\underline{\kappa}_{U}$ lie in the same connected component, and we want to show that $T\cong U$. Recall that two labeled trees with the same number of vertices are isomorphic if and only if for all $1\leq i\leq n$, the set of labels of children of the vertex labeled $i$ are the same in $T$ and $U$, with an analogous statement for the root. Choose a path from $\underline{\kappa}_{T}$ to $\underline{\kappa}_{U}$, call it $\gamma(t)$ ($0\leq t\leq 1$). We can write
@@
so $C_{i}(0)=\underline{C}_{i}^{T}$ and $C_{i}(1)=\underline{C}_{i}^{U}$ for each $1\leq i\leq n$. Note that $\gamma(t)$ is a configuration at every time $t$, so if the $j$th circle is nested in the $i$th circle at some time $t$, this must hold for all times $t$. This shows that for any $i$ and $j$, we have that $\underline{C}_{i}^{T}\vdash\underline{C}_{j}^{T}$ if and only if $\underline{C}_{i}^{U}\vdash\underline{C}_{j}^{U}$. Hence, $i\prec j$ in $T$ if and only if $i\prec j$ in $U$ (with an analogous statement about the root), and so indeed $T\cong U$.
Now suppose $T\cong U$, and we need to prove there is a path in $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$ from $\underline{\kappa}_{T}$ to $\underline{\kappa}_{U}$. We will induct on $n$. If $n=0$ there is nothing to prove, so assume $n\geq 1$. In particular the root is not a leaf. Fix a label $i$ such that the vertex labeled $i$ in $T$, and hence in $U$, is a leaf. Let $\widehat{T}$ be the labeled tree obtained by removing the leaf labeled $i$ from $T$, and relabeling the remaining non-root vertices by replacing $j$ with j1j-1 for all $j>i$. Similarly define $\widehat{U}$. Note that $\widehat{T}\cong\widehat{U}$, so by induction there is a path in $\operatorname{Conf}_{n-1}(S^{1},\mathbb{R}^{2})$ from $\underline{\kappa}_{\widehat{T}}$ to $\underline{\kappa}_{\widehat{U}}$. It is also clear that there is a path in $\operatorname{Conf}_{n-1}(S^{1},\mathbb{R}^{2})$ from $\underline{\kappa}_{\widehat{T}}$ to the labeled configuration obtained from $\underline{\kappa}_{T}$ by deleting the $i$th circle, and similarly from $\underline{\kappa}_{\widehat{U}}$ to $\underline{\kappa}_{U}$ with the $i$th circle deleted. Thus we have a path from $\underline{\kappa}_{T}$ without the $i$th circle to $\underline{\kappa}_{U}$ without the $i$th circle. Let us write
@@
for this path, with $0\leq t\leq 1$, so $C_{j}(0)=\underline{C}_{j}^{T}$ and $C_{j}(1)=\underline{C}_{j}^{U}$ for all $j\neq i$. Note that for each $t$, the path at time $t$ describes a configuration, i.e., for each $t$ the $C_{j}(t)$ are pairwise disjoint for all $j$. Let $k$ be the label on the vertex of $T$ that has the vertex labeled $i$ as a child, and let $j_{1},\dots,j_{p}$ be the labels on the other children of this vertex. For each $t$ let $S(t)$ be the connected open region of the plane obtained by taking the open disk bounded by $C_{k}(t)$ and removing the closed disks bounded by each $C_{j_{\ell}}(t)$ ($1\leq\ell\leq p$). The path $\widehat{\gamma}(t)$ induces, for each $t$, a homeomorphism $h_{t}$ from $S(0)$ to $S(t)$, and these homeomorphisms vary continuously with $t$. Now for our fixed $i$, let $C_{i}^{\prime}(t)$ ($0\leq t\leq 1$) be some choice of a family of continuously varying circles such that $C_{i}^{\prime}(0)=\underline{C}_{i}^{T}$ and for each $t$, $C_{i}^{\prime}(t)$ is contained in the closed disk bounded by $h_{t}(\underline{C}_{i}^{T})$. Write
@@
so our choice of $C_{i}^{\prime}(t)$ ensures that each $\gamma^{\prime}(t)$ is a configuration, and hence this is a path in $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$ from $\gamma^{\prime}(0)=\underline{\kappa}_{T}$ to the configuration
$\kappa^{\prime}\coloneqq(\underline{C}_{1}^{U},\dots,\underline{C}_{i-1}^{U},C_{i}^{\prime}(1),\underline{C}_{i+1}^{U},\dots,\underline{C}_{n}^{U})\text{.}$
Finally, let $\gamma(t)$ be a path obtained by concatenating $\gamma^{\prime}(t)$ with some path from $\kappa^{\prime}$ to $\underline{\kappa}_{U}$ that leaves all but the $i$th circle fixed, and takes the $i$th circle from $C_{i}^{\prime}(1)$ to $\underline{C}_{i}^{U}$. This is possible since $C_{i}^{\prime}(1)$ and $\underline{C}_{i}^{U}$ both lie in the connected region $S(1)$. All in all, we end up with a path $\gamma(t)$ in $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$ from $\gamma(0)=\underline{\kappa}_{T}$ to $\gamma(1)=\underline{\kappa}_{U}$.
Finally, we must prove the last claim. Let $\kappa$ be an arbitrary labeled configuration in $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$, and it suffices to prove that there is a path from $\kappa$ to $\underline{\kappa}_{T}$, where $T\coloneqq T_{\kappa}$. In fact, this works in the same way as the previous part. By induction there is a path between the configurations with the $i$th circles deleted, and then we can extend this to a path between the full configurations since the $i$th circle in $\kappa$ and the $i$th circle in $\underline{\kappa}_{T}$ have the same nesting properties with respect to the other circles in their respective configurations. ∎
@@
Let $T$ and $U$ be trees with $n$ non-root vertices. Then $\underline{\kappa}_{T}$ and $\underline{\kappa}_{U}$ lie in the same connected component of $\operatorname{UConf}_{n}(S^{1},\mathbb{R}^{2})$ if and only if $T\cong U$. Moreover, every connected component of $\operatorname{UConf}_{n}(S^{1},\mathbb{R}^{2})$ contains some $\underline{\kappa}_{T}$ for $T$ a tree, and $T\cong T_{\kappa}$ for each $\kappa$ in this connected component.
@@
First suppose $\underline{\kappa}_{T}$ and $\underline{\kappa}_{U}$ lie in the same connected component, and we need to show that $T\cong U$. Choose some labeling of $T$, and write $\widetilde{T}$ for the resulting labeled tree, so $p(\underline{\kappa}_{\widetilde{T}})=\underline{\kappa}_{T}$. We can lift a path in $\operatorname{UConf}_{n}(S^{1},\mathbb{R}^{2})$ from $\underline{\kappa}_{T}$ to $\underline{\kappa}_{U}$, to a path in $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$ starting at $\underline{\kappa}_{\widetilde{T}}$. The end of this path is some labeled configuration that maps under $p$ to $\underline{\kappa}_{U}$, and so equals $\underline{\kappa}_{\widetilde{U}}$ for some labeled tree $\widetilde{U}$ with underlying tree $U$. By Proposition 3.7, the fact that $\underline{\kappa}_{\widetilde{T}}$ and $\underline{\kappa}_{\widetilde{U}}$ share a connected component implies that $\widetilde{T}\cong\widetilde{U}$, so in particular $T\cong U$.
Now suppose $T\cong U$. Choose a labeling of $T$, and write $\widetilde{T}$ for the resulting labeled tree. Carry the labels on the vertices of $T$ through a choice of isomorphism $T\to U$, to obtain labels on the vertices of $U$, so the resulting labeled tree $\widetilde{U}$ is isomorphic to $\widetilde{T}$ as a labeled tree. By Proposition 3.7, $\underline{\kappa}_{\widetilde{T}}$ and $\underline{\kappa}_{\widetilde{U}}$ share a connected component in $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$. Now the image of a path from $\underline{\kappa}_{\widetilde{T}}$ to $\underline{\kappa}_{\widetilde{U}}$ under the map $p$ is a path in $\operatorname{UConf}_{n}(S^{1},\mathbb{R}^{2})$ from $\underline{\kappa}_{T}$ to $\underline{\kappa}_{U}$.
@@
For $T$ a labeled tree with $n$ non-root vertices, write $\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2})$ for the connected component of $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})$ containing $\underline{\kappa}_{T}$, and call $\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2})$ the space of (labeled) $T$-configurations of $n$ circles in the plane. Similarly, for $T$ an (unlabeled) tree with $n$ non-root vertices, write $\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2})$ for the connected component of $\operatorname{UConf}_{n}(S^{1},\mathbb{R}^{2})$ containing $\underline{\kappa}_{T}$, and call $\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2})$ the space of unlabeled $T$-configurations of $n$ circles in the plane.
@@
Figure 12. Up to isomorphism there are exactly two (unlabeled) trees with $2$ non-root vertices, pictured here. These correspond to unlabeled configurations of $2$ circles in the plane; the first corresponds to the case where the circles are not nested, and the second to the case where the circles are nested.
@@ 4. Braiding automorphism groups of finite trees
Here the proposition is phrased using the “external” viewpoint of semidirect products, i.e., $G=N\rtimes H$ means that $N$ and $H$ are groups, and $H$ acts on $N$ by automorphisms via some implicitly understood (left) action, so the group operation on $N\rtimes H$ is given by $(n,h)(n^{\prime},h^{\prime})=(n(h.n^{\prime}),hh^{\prime})$. In the proof we will actually decompose $\operatorname{Aut}(T)$ using the “internal” viewpoint, i.e., $G=N\rtimes H$ means that $N$ and $H$ are subgroups of $G$, NN is normal, $N\cap H=\{1\}$, and $G=NH$.
@@
Now to prove the decomposition, we will use the internal semidirect product viewpoint. By Lemma 4.1 and its proof, $i$ and $j$ share a block of $\Pi$ if and only if there is an automorphism of $T$ that swaps $T(v)$ and $T(w)$ and fixes everything outside $T(v)\cup T(w)$. These automorphisms generate a copy of $S_{m}^{\Pi}$ inside $\operatorname{Aut}(T)$, since within each block $A\in\Pi$, they generate the relevant copy of $S_{|A|}$. Write $H$ for this subgroup of $\operatorname{Aut}(T)$ isomorphic to $S_{m}^{\Pi}$. Write $N$ for the pointwise stabilizer in $\operatorname{Aut}(T)$ of $\{v_{1},\dots,v_{m}\}$, so $N$ is isomorphic to $\operatorname{Aut}(T(v_{1}))\times\cdots\times\operatorname{Aut}(T(v_{m}))$. Now observe that every element of $\operatorname{Aut}(T)$ can be written as an automorphism in $H$ composed with an automorphism in $N$, so $\operatorname{Aut}(T)=NH$. Moreover, since $S_{m}^{\Pi}$ acts faithfully on $\{v_{1},\dots,v_{m}\}$, we have $N\cap H=\{1\}$. Finally, $N$ is normal, being the kernel of the action on $\{v_{1},\dots,v_{m}\}$, so we conclude that $\operatorname{Aut}(T)=N\rtimes H$. The conjugation action of $H$ on $N$ coincides with the action of $S_{m}^{\Pi}$ on $\operatorname{Aut}(T(v_{1}))\times\cdots\times\operatorname{Aut}(T(v_{m}))$ by permuting entries, so we are done. ∎
@@
This recursive definition makes sense since each $T(v_{i})$ has strictly fewer vertices than $T$. We get that $\operatorname{bAut}(T)$ is built out of iterated semidirect products of subgroups of braid groups. To be clear, in the bottom case when $T$ is the trivial tree, so $m=0$, we have that $B_{0}$ is the trivial group (being generated by the empty set), so for $T$ the trivial tree we have that $\operatorname{bAut}(T)$ is the trivial group. Note that the group $\operatorname{bAut}(T)$ is isomorphic to the group in Lemma 3.1 of [KM22], which lives in the mapping class group of a disk with a puncture for each leaf of $T$, as the stabilizer of a multicurve whose nesting structure is given by $T$ (so the “innermost” circles are viewed as punctures). A special case of this also appears in Section 3 of [CKM]. It would be interesting to pin down this connection explicitly, using a variant of the Birman exact sequence in Theorem 9.1 of [FM12].
@@
As for the pure case, it is clear from Definition 4.5 that $\operatorname{pbAut}(T)\cong PB_{3}\times PB_{3}\times PB_{2}$ for this $T$, since two vertices have $3$ children, one vertex has $2$ children, and all other vertices have $1$ or 0 children.
@@
If every non-root vertex of $T$ is a child of the root, and there are $n$ non-root vertices, then $\pi_{1}(\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2}))\cong\operatorname{bAut}(T)$ by virtue of both being isomorphic to $B_{n}$.
@@
First note that $\operatorname{bAut}(T)\cong B_{n}$ in this case by construction. Now for $C$ a circle in the plane, let $D_{C}$ be the disk in the plane obtained by taking the union of $C$ with the bounded component of its complement. Since every non-root vertex of $T$ is a child of the root, every configuration $\{C_{1},\dots,C_{n}\}$ in $\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2})$ has the property that no $C_{j}$ is nested in any $C_{i}$. Thus, we have a well defined map $\{C_{1},\dots,C_{n}\}\mapsto\{D_{C_{1}},\dots,D_{C_{n}}\}$ from $\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2})$ to $\operatorname{UConf}_{n}(D^{2},\mathbb{R}^{2})$. This map is clearly a homeomorphism, so $\pi_{1}(\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2}))\cong B_{n}$ by Lemma 2.6. ∎
An analogous procedure shows that $\pi_{1}(\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2}))\cong\operatorname{pbAut}(T)$ for labeled trees $T$ of this form, thanks to both being isomorphic to $PB_{n}$.